### v2

# Landau-Lifshitz damping from Lindbladian dissipation in quantum magnets

Abstract
As of now, the phenomenological classical Landau-Lifshitz (LL) damping of magnetic order is not conceptually linked to the quantum theory of dissipation of the Lindbladian formalism which is unsatisfactory for the booming research on magnetic dynamics. Here, it is shown that LL dynamics can be systematically derived from Lindbladian dynamics in a local mean-field theory for weak external fields. The derivation also extends the LL dynamics beyond the orientation $\vec{m}/|\vec{m}|$ to the length $|\vec{m}|$ of the magnetization. A key assumption is that the Lindbladian dissipation adapts to the non-equilibrium $H(t)$ instantaneously to lower its expectation value.
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
Magnetic dynamics are a topic attracting an enormous interest these days because it is crucial for making progress in magnonics [1], i.e., for the general idea to use magnetic degrees of freedom for information storage and handling. We stress that this requires to deal with non-equilibrium physics. Ferromagnetic [2, 3] and antiferromagnetic systems [4, 5] as well as spin textures [6] are considered with their respective advantages and disadvantages. For substantial progress numerous experimental challenges have to be overcome and well-founded theoretical understanding is indispensable. This requires to capture the relevant Hamiltonian dynamics as well as the relaxation processes. While the former is often known from experimental measurements or from ab initio calculations, the latter is generically described phenomenologically on a classcial level by the Landau-Lifshitz (LL) equation or the Landau-Lifshitz-Gilbert (LLG) equation [7, 8, 9]. These equations, though phenomenological, are fundamental to realistic descriptions of magnetism away from equilibrium [10].
Since spins as the elementary objects of magnetism are quantum entities and in view of the importance of the LL(G) equations it is not surprising that there are intensive efforts to establish a quantum foundation for them [11, 12, 13, 14, 15, 16] uncovering similarities between Lindbladian dynamics for magnons, i.e., usual bosons, and results from the LLG equations. Considering magnons as damped harmonic oscillators, it has been indeed possible to identify a regime where the Lindblad and the classical LLG equations coincide in linear order of the deviations from equilibrium if a particular bath structure is assumed [15]. Yet, we would like to address the issue more generally, i.e., independent of the bath properties as long as the bath allows for a Lindblad description and independent of a small-angle approximation. Furthermore, it is pointed out in Ref. [16] that for an LL(G) description to be valid the non-local entanglement between the spins needs to vanish since the LL(G) behavior is classical. Using a sophisticated Lindblad approach for systems of four spins it is concluded that an LL(G) description may only be valid for $S>1$, i.e., for large spin. Yet, very small systems cannot mimic the behavior of large spin systems in higher dimensions. It is known, that the behavior of spins in ordered quantum magnets in higher dimensions can be captured well in mean-field approaches because the effect of quantum fluctuations is reduced. We assume such a high-dimensional system where a local mean-field approach can be used neglecting non-local entanglement. The derivation based on such a local mean-field approach is mathematically rigorous. We stress that a local mean-field approach still captures varying lengths of the spin expectation values, i.e., it goes beyond a purely classical treatment.
Another idea is to establish a quantum mechanical equation that results in the established classical equations in the classical limit [11, 12, 14]. The assumptions necessary to achieve this goal are a strong non-Hermitian Hamiltonian [11, 12, 17] or a self-referential term in the von-Neumann equation where the density matrix appears twice [14]. The appearance of a non-Hermitian part has been justified for Bose-Einstein condensates on heuristic grounds in view of conservation laws [18, 19, 20]. Further studies are called for justifying these assumptions on the basis of physical processes.
In parallel, a number of approaches exists which compute the Gilbert damping parameter obtaining reasonable values on the basis of time-dependent or non-equilibrium quantum mechanical approaches addressing the baths explicitly [21, 22, 23, 24, 25, 26, 27, 13, 28, 29]. The resulting master equations, however, do not always amount up to Lindbladian dynamics [30]. Still, a quantum mechanical treatment of the baths must lead to Lindbladian dynamics if its preconditions are met: weak-coupling between system and bath, justification of Born-Markov approximation (bath correlations decay faster than the system’s internal dynamics), and the rotating wave approximation [31]. Thus, it must be possible to derive the damping in the LL(G) equations from standard quantum mechanics for open quantum systems in a regime where both descriptions of dissipation are justified. It is the objective of this paper to provide such a systematically controlled derivation. It is not our goal to describe damping beyond the validity of a Lindblad approach nor do we aim at incorporating non-local entanglement to atomistic spin simulations [10, 32, 33, 34] because we rely on local mean-field theory. But we stress that changes of the spin length as computed from the squares of the linear spin expectation values are captured by local mean-field theories.
The Lindblad formalism is a well-established and successful theory for relaxation of quantum systems under certain conditions. The key foundation is to consider an open quantum system coupled to a bath [31] leading to the Lindblad master equation including a so-called dissipator. If the damping of a magnetic systems can be described by the LL(G) equations, which is governed by a single relaxation rate, the latter should be derivable from Lindbladian dynamics. So far, however, Lindblad dynamics and LL(G) dynamics are not linked by a mathematically rigorous derivation, except that for special systems their outcomes are the same [15]. This situation is unsatisfactory since the spins are quantum objects and cannot have their own relaxation independent from the general quantum theory. In this study, we will fill this gap and link Lindbladian dynamics and LL(G) dynamics under certain mild conditions. We show that the latter results from the former in a general limit based on a local mean-field theory, i.e., without non-local correlations [16]. This is justified in higher dimensions. We identify additional terms in the systematic derivation which describe the change of the length of the magnetization, i.e., $|\vec{m}|$.
2 Brief Recapitulation of Landau-Lifshitz and Lindblad Formalisms
As a brief recapitulation, the LL equation reads
$\frac{d\hat{\vec{m}}}{dt}=\hat{\vec{m}}\times\vec{h}_{0}-\lambda\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})$ (1a)
if the Hamilton function reads $H=-\vec{h}_{0}\cdot\hat{\vec{m}}$ and $\hat{\vec{m}}$ is the normalized magnetization of a spin and $\vec{h}_{0}$ a fixed external magnetic field with $h_{0}=g\mu_{\text{B}}B$. The LLG equation reads
$\frac{d\hat{\vec{m}}}{dt}=\hat{\vec{m}}\times\vec{h}_{0}-\lambda\hat{\vec{m}}\times\frac{d\hat{\vec{m}}}{dt}.$ (1b)
Both equations are equivalent within ${\cal O}(\lambda^{2})$ [35, 12] because (1b) can be transformed to
$\frac{d\hat{\vec{m}}}{dt}=\frac{1}{1+\lambda^{2}}\left[\hat{\vec{m}}\times\vec{h}_{0}-\lambda\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})\right].$ (2)
Inserting (1b) in itself and using the bac-cab rule $a\times(b\times c)=b(a\cdot c)-c(a\cdot b)$ for $a,b,c\in\mathds{R}^{3}$ as well as $\hat{\vec{m}}\cdot\frac{d\hat{\vec{m}}}{dt}=0$ leads to (2). Hence the LL and LLG equation stand for almost the same dynamics except for a renormalization of the time $t\to(1+\lambda^{2})t$. The difference only appears in second order in $\lambda$.
Lindbladian dynamics for an operator $A$ in the Heisenberg picture is given by the adjoint Master equation which reads at zero temperature [31, 36]
$\frac{d\langle A\rangle}{dt}=i\langle[H,A]\rangle+\frac{1}{2}\sum_{l}\gamma_{l}\Big{\langle}[B_{l},A]B^{\dagger}_{l}+B_{l}[A,B^{\dagger}_{l}]\Big{\rangle}$ (3)
where the $\gamma_{l}$ are the decay rates of the various dissipative channels which are defined by the Lindblad operators $B_{l}$. It is implied that $B_{l}$ increments the energy in $H$ by $\hbar\omega_{l}>0$ while $B_{l}^{\dagger}$ decreases the energy by the same amount. At finite temperatures, an additional term occurs which is derived and discussed in App. D. Finite temperature does not change the results qualitatively, but increases relaxation and prevents that the magnetizations attains its maximum values.
3 Single Spin in Time-Dependent Field in Lindblad Formalism
In a first step, we consider a single spin coupled to a static magnetic field which we assume for simplicity to point into $z$ direction $\vec{h}_{0}=h_{0}(0,0,1)^{\top})$. The minimal relaxation channel is given by $B=S^{-}$ with rate $\gamma$. Equation (3) implies
$\frac{d\langle S^{z}\rangle}{dt}$ $=2\gamma S(S-\langle S^{z}\rangle)$ (4a)
$\frac{d\langle S^{+}\rangle}{dt}$ $=-(ih_{0}+\gamma S)\langle S^{+}\rangle$ (4b)
for the spin $S$; with appropriate definitions for longitudinal and transversal relaxation these are the Bloch equations [37]. Qualitatively, they also hold at finite temperatures, see App. D, where also the absorption process induced by $B^{\dagger}$ contributes leading to terms proportional to the bosonic occupation. The equations (4) are exact for $S=1/2$; for general spin they are valid if the spin state deviates only little from the state in the polarized $z$ direction, i.e., if only the magnetic quantum numbers $S$ and $S-1$ occur. We will discuss the justification of this approximation below. The solutions of (4) read $\langle S^{z}\rangle(t)=(\langle S^{z}\rangle(0)-S)\exp(-2\gamma St)+S$ and $\langle S^{+}\rangle(t)=\exp(-(ih_{0}+\gamma S)t)\langle S^{+}\rangle(0)$. This reflects standard longitudinal relaxation combined with damped transverse precession, i.e., the physically expected result. Yet, this behavior is strikingly different from what follows from Eqs. (1)! Yet this is not the end of the story.
As an important intermediate step we consider the magnetic field $\vec{h}$ to be time dependent. For the Lindbladian approach to be valid, the bath dynamics needs to be faster than the relaxation and the internal dynamics (Born-Markov approximation) [31]. In generic solid state systems, the phonons constitute the bath and their characteristic energy scale is set by the Debye frequency $\omega_{\rm Debye}$. Hence, we assume that $\hbar\omega_{\rm Debye}$ is larger than the magnetic coupling $J$. But it is not necessary to specify the physical origin of the bath except that it displays a rapid internal dynamics which allows us to assume that the bath can adapt to $\vec{h}(t)$ at each instant of time quickly, i.e., the Lindblad operator $B(t)$ adapts to the instantaneous magnetic field. Hence the direction of relaxation is given by $\vec{h}(t)$, i.e., towards aligning the spin with its momentary effective field. In App. A we carry out the systematic rotation of Eqs. (4) and we obtain for $\vec{m}\coloneqq\langle\vec{S}\rangle$
$\frac{d\vec{m}}{dt}=\vec{m}\times\vec{h}+2\gamma S(S\vec{n}-\vec{m})-\gamma S\vec{n}\times(\vec{n}\times\vec{m}),$ (5)
where $\vec{n}=\vec{h}/h$ is the time-dependent unit vector in the direction of the magnetic field. In the first term on the right hand side one recognizes the precession and the last term is a double cross product, but twice with $\vec{n}$ instead of $\hat{\vec{m}}$. The self-reference is missing which is present in Eqs. (1).
4 Local Mean-Field Approach and its Numerical Evaluation
Here, the next key element enters. We do not attribute the self-reference to an assumed modification of the Schrödinger equation, but to a local mean-field approximation considering the ensemble of spins. For simplicity, we treat a translational invariant ferromagnet in which each spin is subject to the external field $\vec{h}_{0}=h_{0}(0,0,1)^{\top})$ and to the internal exchange field generated by the surrounding spins interacting with the spin under study so that
$\vec{h}=\vec{h}_{0}+J\vec{m},$ (6)
where $J$ represents the sum over all exchange couplings acting on the considered spin. Note that this approach is easily generalized to arbitrary spin textures defining a local effective magnetic field $\vec{h}_{i}\coloneqq\sum_{j}J_{ij}\vec{m}_{j}$. Eventually, one obtains LL equations for each site. If each spin interacts with a sufficiently large number of other spins or if the spin orientation only changes slowly in space coarse graining can be performed as well to introduce spin densities. Thus, a number of extensions is clearly possible.
Figure 1: Solutions (dashed lines) of the Landau-Lifshitz equation (1a) for $\lambda=\gamma/J=0.2$ and $S=1/2$. Solutions (solid lines) of the Lindbladian equation (5). The initial direction of the magnetization of length $S$ is tilted by $\pi-0.1$ with respect to the $z$ direction.
Inserting (6) into (5) yields a closed, self-referential set of equations for the magnetization $\vec{m}$. Still, it seems to be distinctly different from Eqs. (1), but this is illusive. A first step to see this is to estimate orders of magnitude. Generic ferromagnets have a Curie temperature of about $10^{3}$ K corresponding roughly to an internal magnetic field of $10^{3}$ T. Externally applied fields range from 1 mT to 10 T. Thus, we face the situation that $h_{0}$ is about (and even more than) 1000 times weaker than the internal field $JS$. This suggests that one must focus on the limit of weak external fields, i.e., on $h_{0}\ll J$.
Figure 1 confirms the conjecture that the LL dynamics is the weak-field limit of the Lindbladian dynamics. The dashed lines show the solutions of Eq. (1a) while the solid lines in the same color represent the solutions of Eq. (5) using (6). The time axis is displayed in units of $1/h_{0}$ to show the Larmor precessions clearly. The relaxation rate $\gamma$ is kept constant relative to the exchange $J$, i.e., the parameter $\lambda=\gamma/J$ is fixed.
Clearly, far in the weak field limit for $J/h_{0}=100$ the agreement is excellent; only a minor shift in the relaxation of the $z$ component is discernible, see panel (a). Lowering the ratio $J/h_{0}$ the deviations increase, but the qualitative behavior still is close to the LL dynamics, see panel (b). Figure 1(c) shows that for significant lower ratios the qualitative (!) behavior changes: the $z$ component essentially displays exponential relaxation as one would expect from (4). These numerical observations corroborate the above ypothesis that the LL dynamics results from a Lindblad equations in the limit of weak external fields $h_{0}$ in comparison to the strong internal effective fields $J$. If this limit is reached by increasing $J$ while keeping $h_{0}$ constant, it is important to increase the relaxation rate $\gamma$ as well such that $\lambda=\gamma/J$ remains constant.
At this point, we emphasize that the use of Eq. (6) for large ratios $J/h_{0}$ implies that the spin orientation is always very close to the direction of the effective magnetic field $\vec{h}$, not to $\vec{h}_{0}$. This justifies the use of the approximation necessary to establish (4) for $S>1/2$. For $S=1/2$, the differential equations (4) are exact anyway.
5 Analytical Derivation of Generalized Landau-Lifshitz Equation
One could use more numerical data for supporting the above hypothesis. But it is by far more general to back the numerical agreement by an analytical argument which captures the dependencies on all parameter. If the LL dynamics is the weak-field limit of the Lindblad dynamics the LL equation must be derivable by a systematic expansion in the external field $h_{0}$. Expanding the right hand side of (5) in linear order in $h_{0}$ using $\vec{n}=\vec{h}/h$ and (6) yields
$\frac{1}{S}\frac{d\vec{m}}{dt}=\frac{\vec{m}}{S}\times\vec{h}_{0}+2\gamma S(\hat{\vec{m}}-\vec{m}/S)-\frac{\gamma}{J}C\,\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})$ (7)
with $C\coloneqq 2S/|\vec{m}|-1$. the mathematical details of the expansion are included in App. B. We distinguish here between the magnetization with arbitrary length $\vec{m}$ and the normalized one $\hat{\vec{m}}$. Inspecting the three terms on the right hand side of (7) the first one induces the precession. It is often quite fast. The second one is also a fast one; it could even be faster than the precession depending on the applied field. Clearly, it is faster than the third one since we are considering the regime of large $J$. The second term ensures that the magnetization converges to its saturation value $m_{\infty}=S$ at zero temperature quickly on the time scale $1/\gamma$. At finite temperature, the saturation value $m_{\infty}$ is lower, see the derivation and discussion in App. D.
The second term in Eq. (7) represents a systematic extension of the LL equations for non-saturated magnetizations still fulfilling all required symmetries. Clearly, it goes beyond a purely classical dynamics because the vector length $|\vec{m}|$ will vary as well; it is a consequence of the local mean-field approach. If, however, we start from an initial magnetization that takes its saturation value this term has no effect anymore and $|\vec{m}|=m_{\infty}=S$ holds for all subsequent times. This implies $C=1$ and that the time derivative on the left hand side equals $\frac{d\hat{\vec{m}}}{dt}$ as well as the precession term equals $\hat{\vec{m}}\times\vec{h}_{0}$. The third term on the right hand side represents the LL damping term if we set $\lambda=\gamma/J$. This analytically supports the hypothesis that the LL dynamics results from the Lindblad dynamics in the limit of weak external fields. It is remarkable that in the case of saturation $|\vec{m}|=S$ the time scale $1/\gamma$ disappears completely and only the time scales $1/h_{0}$ and $J/(\gamma h_{0})$ survive. The fact that for a magnetization almost parallel to the external field the change $\frac{d\vec{m}}{dt}$ becomes very small, even vanishing for parallel orientation, is an essential feature of the LL equation. Here it follows automatically from the Lindblad formalism.
Thus, we eventually derived the LL equation (1a) from Lindbladian dynamics and identified the dimensionless damping parameter $\lambda$ by the ratio of the relaxation rate $\gamma$ to the internal energy scale $J$. The latter should be larger than $\gamma$ in order that the weak-coupling assumption in deriving Lindbladian dynamics hold [31]. Hence we can safely assume $\lambda\leq 1$ so that no unphysical behavior results from the LL equation. We cannot derive the LL equation beyond the realm of Lindbladian dynamics. In addition, we studied larger values of $\lambda$ to see whether the LL equation (1a) or the LLG equation (1b) corresponds better to the Lindbladian dynamics (5), see App. C. In accordance with the analytical derivation of the weak-field limit we find that the LL equation represents the correct limit of weak external fields best. We point out, however, that the LL equation is known to display unphysical behavior for $\lambda>1$ so that restrict ourselves to $\lambda<1$.
It is interesting that for $\vec{m}\neq m_{\infty}\hat{\vec{m}}$ Eq. 7 extends the usual LL equation (1a) systematically, see also App. D. This can be used to raise simulations from a classical LL level to the quantum mechanical level of local mean fields although non-local entanglement is still neglected.
6 Conclusions and Outlook
In summary, up to now there was a conceptual gap between the established Lindbladian theory for dissipation in open quantum systems and the phenomenological theory of Landau, Lifshitz and Gilbert for the damping of the temporal behavior of magnetic order. In view of the humongous interest in magnetic dynamics this represented a flaw. So far, the differences of the two approaches were emphasized, for instance the change of entropy in Lindblad treatments, see e.g. Ref. [38], and the conservation of entropy in the assumed quantum analogues of the LL(G) equations [11, 14]. In the present study, we filled the conceptual gap and reconciled the fundamental quantum description in terms of a Lindbladian master equation and the phenomenological classical LL(G) description. In addition, a systematic extension for the evolution of the length of the expectation value of $\vec{S}$ has been found. This underlines that the Lindblad approach, though based on a local mean-field theory, reaches beyond a merely classical description.
Three key points constitute the important ingredients: (i) the Lindbladian relaxation is constantly adapted to the instantaneous Hamilton operator and its energies, i.e., the relaxation always acts such as to minimize the instantaneous energy; (ii) the self-reference in the LL(G) equations is explained not by ad hoc modifications of the Schrödinger equation, but by a local mean-field approximation; (iii) finally the internal energy scales of the quantum system need to be much larger than the external ones, here $|J|\gg|h_{0}|$. Key point (i) requires that the internal bath dynamics is faster than the system’s dynamics. This is a standard assumption in the derivation of the derivation of the Lindblad master equation. These ideas allowed us to derive the Landau-Lifshitz equation analytically from Lindbladian dynamics and to confirm this by numerical calculations. We did not observe a trace of Landau-Lifshitz-Gilbert dynamics for large values of $\lambda=\gamma/J$.
The possibility to derive LL dynamics from Lindbladian dynamics provides the general context for computing damping parameters from elaborate theories describing the reservoirs explicitly which are responsible for the relaxation of quantum system [21, 22, 23, 24, 25, 26, 27, 13, 28]. The above ideas can also be applied directly to quantum magnets beyond ferromagnets as along as they are treated by a local mean-field theory. An important example are quantum antiferromagnets gaining more and more interest for the prospect of faster data handling [4, 39] or general spin textures [6]. We stress that quantum fluctuations of the spin ensemble itself are not captured; they would require advanced mean-field techniques [40, 41] or modified approaches to atomistic spin dynamics [10, 32, 33, 34].
Another extension can consist in using Eq. (7) instead of the LL equation in atomistic spin simulations [9, 42, 43, 10]. Such approaches are also employed to describe systems of Josephson junctions [44, 45, 46]. This would allow for taking changes in the spin length into account, promoting the classical calculation to a local mean-field one if the external fields are weak with respect to the internal ones. If the weak-field limit is not justified one can equally well resort to Eq. (5) which is based on a local mean-field approach, but does not require a particular energy hierarchy between internal and external magnetic energies. For including thermal fluctuations by stochastical noise terms, the conditions for reaching equilibrium need to be verified [43].
Recently, high frequencies in the spin dynamics of ferromagnets have been predicted and experimentally identified which are attributed to spin inertia and a resulting nutation term in the Landau-Lifshitz-Gilbert equations [47, 48, 49, 50]. Mathematically, a second time derivative of the magnetization is included. This poses the question whether such a term can result from a Lindblad formalism by extending the above reasoning. Alternatively, one can attempt to understand the observed signatures by Eqs. (5) or (7) directly
The success of the presented calculations clearly suggests that Lindbladian dissipation captures the physics of relaxation very well. The above key idea (i) that it must be constantly adapted to the instantaneous Hamiltonian in time-dependent, non-equilibrium situations is generally applicable if the energy scales are such that the internal bath dynamics are the fastest. Thus, the reported findings may pave the way to studies of relaxation in a plethora of non-equilibrium quantum systems.
Acknowledgements
The author is thankful to Timo Gräßer, Asliddin Khudoyberdiev, and Joachim Stolze for useful discussions.
Funding information
This research has been supported by the Deutsche Forschungsgemeinschaft (DFG) in project UH 90-14/1 and by the Stiftung Mercator in project KO-2021-0027
Appendix A Derivation of Lindbladian Dynamics for Time-Dependent Magnetic Field
The aim is to derive Eq. (5). This implies that we carry out a rotation in spin space to account for the magnetic field varying in time. Concomitantly, the Lindblad operator varies in time as well since it is assumed to relax the spin at each instant of time according to the instantaneous direction of the field, i.e., such that the spin approaches parallel orientation to the magnetic field.
We choose $\vec{n}\coloneqq\vec{h}/h$ and an orthogonal unit vectors $\vec{p}$ so that they for a orthonormal basis of $\mathds{R}^{3}$ together with $\vec{q}\coloneqq\vec{n}\times\vec{p}$. Their completeness is expressed by
$\vec{n}\vec{n}^{\top}+\vec{p}\vec{p}^{\top}+\vec{q}\vec{q}^{\top}=\mathds{1}_{3}$ (A.1)
where $\mathds{1}_{3}$ is the $3\times 3$ identity matrix. Differentiation of the completeness with respect to time yields
$\frac{d\vec{n}}{dt}\vec{n}^{\top}+\vec{n}\frac{d\vec{n}}{dt}^{\top}+\frac{d\vec{p}}{dt}\vec{p}^{\top}+\vec{p}\frac{d\vec{p}}{dt}^{\top}+\frac{d\vec{q}}{dt}\vec{q}^{\top}+\vec{q}\frac{d\vec{q}}{dt}^{\top}=0$ (A.2)
which we will use below. For a concise notation we introduce the complex vector
$\vec{z}\coloneqq\vec{p}+i\vec{q}\in\mathds{C}^{3}.$ (A.3)
Then we can compute the longitudinal and transverse spin expectation value using $\vec{m}=\langle\vec{S}\rangle$
$S_{\parallel}$ $\coloneqq\vec{n}\cdot\vec{m}$ (A.4a)
$S_{+}$ $\coloneqq\vec{z}\cdot\vec{m}.$ (A.4b)
These values acquire their time dependence from Eqs. (4) and from the time dependence of $\vec{n}$ and $\vec{z}$
$\frac{dS_{\parallel}}{dt}$ $=\gamma 2S(S-S_{\parallel})+\frac{d\vec{n}}{dt}\cdot\vec{m}$ (A.5a)
$\frac{dS_{+}}{dt}$ $=-(ih+\gamma S)S_{+}+\frac{d\vec{z}}{dt}\cdot\vec{m}.$ (A.5b)
Reconstructing $\vec{m}$ from the two components defined in (A.4) yields
$\vec{m}$ $=\vec{n}S_{\parallel}+\frac{1}{2}\vec{p}\,(S_{+}+S_{+}^{*})-\frac{i}{2}\vec{q}\,(S_{+}-S_{+}^{*})$ (A.6a)
$=\vec{n}S_{\parallel}+\Re(\vec{z}^{\,*}S_{+}).$ (A.6b)
Computing the derivative of $\vec{m}$ from (A.6b) and (A.5) yields
$\frac{d\vec{m}}{dt}$ $=\frac{d\vec{n}}{dt}(\vec{n}\cdot\vec{m})+\vec{n}\Big{(}\frac{d\vec{n}}{dt}\cdot\vec{m}\Big{)}$ (A.7a)
$\quad+\Re\Big{(}\frac{d\vec{z}^{\,*}}{dt}(\vec{z}\cdot\vec{m})+\vec{z}^{\,*}(\frac{d\vec{z}}{dt}\cdot\vec{m})\Big{)}$ (A.7b)
$+\gamma 2S\vec{n}(S-\vec{n}\cdot\vec{m})-\Re\big{(}(ih+\gamma S)\vec{z}^{\,*}(\vec{z}\cdot\vec{m})\big{)}.$ (A.7c)
The lines (A.7a) and (A.7b) cancel due to (A.2). Using
$\Re(\vec{z}^{\,*}\vec{z}^{\top})=\vec{p}\vec{p}^{\top}+\vec{q}\vec{q}^{\top}=\mathds{1}_{3}-\vec{n}\vec{n}^{\top},$ (A.8)
where the second equation results from (A.1), as well as
$\Im(\vec{z}^{\,*}(\vec{z}\cdot\vec{m}))$ $=\vec{p}(\vec{q}\cdot\vec{m})-\vec{q}(\vec{p}\cdot\vec{m})$ (A.9a)
$=\vec{m}\times(\vec{p}\times\vec{q})=\vec{m}\times\vec{n}$ (A.9b)
leads us to
$\frac{d\vec{m}}{dt}=\vec{m}\times\vec{h}+\gamma 2S^{2}\vec{n}-\gamma S\vec{n}(\vec{n}\cdot\vec{m})-\gamma S\vec{m}.$ (A.10)
To the right hand side we add and subtract $\gamma S\vec{m}$ so that we can use $\vec{m}-\vec{n}(\vec{n}\cdot\vec{m})=-\vec{n}\times(\vec{n}\times\vec{m})$ to obtain
$\frac{d\vec{m}}{dt}=\vec{m}\times\vec{h}+\gamma 2S(S\vec{n}-\vec{m})-\gamma S\vec{n}\times(\vec{n}\times\vec{m}).$ (A.11)
This is the sought equation (5).
Appendix B Linear Order in the External Field $h_{0}$
Here we carry out an expansion of Eq. (5), identical to (A.11), in orders of the external magnetic field $h_{0}$. There is no zeroth order so that we aim at the linear order. The first term in (A.11) takes the form
$\vec{m}\times\vec{h}=\vec{m}\times\vec{h}_{0}$ (B.1)
because $\vec{m}\times\vec{m}=0$; this is linear in $h_{0}$. In the second term, we have to use
$h^{2}=J^{2}|\vec{m}|^{2}+2J\vec{h}_{0}\cdot\vec{m}+h_{0}^{2}$ (B.2)
to reach
$\vec{n}$ $=\frac{\vec{h}}{h}$ (B.3a)
$=\frac{\vec{h}_{0}}{J|\vec{m}|}+\hat{\vec{m}}\Big{(}1-\frac{\vec{h}_{0}\cdot\hat{\vec{m}}}{J|\vec{m}|}\Big{)}+{\cal O}(h_{0}^{2})$ (B.3b)
$=\hat{\vec{m}}+\frac{1}{J|\vec{m}|}(\vec{h}_{0}-\hat{\vec{m}}(\hat{\vec{m}}\cdot\vec{h}_{0}))+{\cal O}(h_{0}^{2})$ (B.3c)
$=\hat{\vec{m}}-\frac{1}{J|\vec{m}|}\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})+{\cal O}(h_{0}^{2})$ (B.3d)
where we use the normalized magnetization $\hat{\vec{m}}=\vec{m}/|\vec{m}|$. The third term in (A.11) simplifies according to
$\vec{n}\times(\vec{n}\times\vec{m})$ $=\frac{\vec{h}\times(\vec{h}\times\vec{m})}{h^{2}}$ (B.4a)
$=\frac{\vec{h}\times(\vec{h}_{0}\times\vec{m})}{h^{2}}$ (B.4b)
$=\frac{J\vec{m}\times(\vec{h}_{0}\times\vec{m})}{J^{2}|\vec{m}|^{2}}+{\cal O}(h_{0}^{2})$ (B.4c)
$=-\frac{\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})}{J}+{\cal O}(h_{0}^{2}).$ (B.4d)
Adding the results (B.1), (B.3d), and (B.4d) with the appropriate prefactors yields
$\frac{d\vec{m}}{dt}=\vec{m}\times\vec{h}_{0}+\gamma 2S(S\hat{\vec{m}}-\vec{m})-\frac{\gamma S}{J}\Big{(}2\frac{S}{|\vec{m}|}-1\Big{)}\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0}).$ (B.5)
Dividing by $S$ yields the claimed equation (7).
Appendix C Effect of Large Landau-Lifshitz Damping
The figures in the article and the analytical calculation presented above in the previous section indicate that the LL equation (1a) captures the limit of weak external fields of the Lindbladian dynamics best. Yet, the damping value used in the numerical computation, $\lambda=0.2$, is fairly small and any possible effect resulting from the LLG equation (1b) on the dynamics would be only $4\%$, see Eq. (2). Hence, we study a larger values of $\lambda$. Yet we stick to values $\lambda<1$ since it is known that the LL equation (1a) yields unphysical behavior for $\lambda>1$. Moreover, we only need to identify possible quadratic effects which distinguish Eq. (1a) from Eq. (2). Thus, to corroborate the observation that the Lindbladian dynamics reduces to LL dynamics rather than to LLG dynamics we study the case $\lambda=0.5$ in Fig. 2.
Figure 2: Solutions (dashed lines) of the Landau-Lifshitz equation (1a) for $\lambda=\gamma/J=0.2$ and $S=1/2$. Solutions (solid lines) of the Lindbladian equation (5). The initial direction of the magnetization of length $S$ is tilted by $\pi-0.1$ with respect to the $z$ direction.
Of course, the damping takes place faster by a factor of $2.5$ than in Fig. 1. In summary, however, we find again that the agreement for large $J$ in panel (a) is almost perfect while deviations become discernible for intermediate $J$ in panel (b). But the qualitative behavior remains the same. We do not observe any shift in the oscillations of the $S^{x}$ component relative to the oscillations in the corresponding LL curve. Hence, we do not see any sign of a renormalization of the precession term. From the LLG equation (2), one would have expected a $25\%$ effect. Also the dynamics of the $S^{z}$ remains very much the same: it compares similarly to the LL dynamics as we observed in Fig. 1. Definitely, there is no $25\%$ effect. Hence, the weak-field limit is given by the LL dynamics (1a) rather than by the LLG dynamics (1b).
Appendix D Effect of Finite Temperature on the Dissipation
At finite temperature, the dissipator denoted in Eq. (3) must be modified [36] by adding the term
$T_{1}\coloneqq\frac{1}{2}\sum_{l}\gamma_{l}n_{\text{B}}(\Delta_{l})\Big{\langle}\big{[}[B_{l},A]B^{\dagger}_{l}\big{]}+\big{[}B_{l}[A,B^{\dagger}_{l}]\big{]}\Big{\rangle}$ (D.1)
where $n_{\text{B}}$ is the bosonic occupation number at inverse temperature $\beta$ given by $n_{\text{B}}=(\exp(\beta\Delta_{l})-1)^{-1}$ as usual. We stress that $T_{1}$ includes the effects of the Lindblad operators $B_{l}$ inducing decay as well as the effects of the Lindblad operators $B_{l}^{\dagger}$ inducing absorption processes. Here, $\Delta_{l}$ is the energy increment induced by the Lindblad operator $B_{l}$. Specifying to the considered case of a spin in a magnetic field, we only have one $l$ and $B=S^{-}$ with $\Delta=J|\vec{m}|$. Carrying out the calculation we find the extended version of Eq. (4)
$\frac{d\langle S^{z}\rangle}{dt}$ $=2\gamma S(S-(1+n_{\text{B}}/S)\langle S^{z}\rangle)$ (D.2a)
$\frac{d\langle S^{+}\rangle}{dt}$ $=-(ih_{0}+\gamma S(1+n_{\text{B}}/S))\langle S^{+}\rangle.$ (D.2b)
To account for the changes, we define $\widetilde{\gamma}\coloneqq\gamma(1+n_{\text{B}}/S)$ allowing us to rewrite the above equations as
$\frac{d\langle S^{z}\rangle}{dt}$ $=2\widetilde{\gamma}S(S-\langle S^{z}\rangle)-2\widetilde{\gamma}S\delta$ (D.3a)
$\frac{d\langle S^{+}\rangle}{dt}$ $=-(ih_{0}+\widetilde{\gamma}S)\langle S^{+}\rangle$ (D.3b)
with $\delta\coloneqq n_{\text{B}}/(1+n_{\text{B}}/S)$. These changes modify Eq. (5) to
$\frac{d\vec{m}}{dt}=\vec{m}\times\vec{h}+2\widetilde{\gamma}S((S-\delta)\vec{n}-\vec{m})-\widetilde{\gamma}S\vec{n}\times(\vec{n}\times\vec{m})$ (D.4)
and in linear order in $h_{0}$ Eq. (7) becomes
$\frac{1}{S}\frac{d\vec{m}}{dt}$ $=\frac{\vec{m}}{S}\times\vec{h}_{0}+2\widetilde{\gamma}S\left(\frac{\hat{\vec{m}}}{1+n_{\text{B}}/S}-\frac{\vec{m}}{S}\right)$
$\quad-\frac{\widetilde{\gamma}}{J}C_{T}\,\hat{\vec{m}}\times(\hat{\vec{m}}\times\vec{h}_{0})$ (D.5)
with $C_{T}\coloneqq 2S/[|\vec{m}|(1+n_{\text{B}}/S)]-1$. Note that the length $m_{\infty}$ of the magnetization in the stationary regime is reduced from $S$ at zero temperature to $S/(1+n_{\text{B}}/S)<S$ at finite temperature. For $S=1/2$ this is the same result as in a local mean-field theory. For larger values of $S$ only the leading influence of $T>0$ is captured because Eq. (4) is not exact for $S>1/2$, but only a good approximation close to full polarization. Of course, this can be remedied if required, for instance by modifications in an atomistic spin simulation [10, 32, 34].
Inspecting the equations (D.4) and (D) the main difference relative to the dynamics at zero temperature is an enhanced relaxation rate $\widetilde{\gamma}>\gamma$ and a difference in the long-time limit $m_{\infty}$ of the absolute value of the magnetization $\vec{m}$. As expected, finite temperature enhances the relaxation and reduces $m_{\infty}$ from its value at zero temperature. Hence, if one starts from the long-time limit, which is also the equilibrium value in mean-field theory for $S=1/2$, the equations are the same as at zero temperature except for the change in the relaxation rate. In particular, we find that the Landau-Lifshitz equation can be deduced from the Lindbladian under mild and plausible assumptions. This extends the key result of the main text to finite temperatures.